OpenStudy (swag):

Can someone show me step by step how to solve this completely: 5(2x - 7) = 33 - 3(6x + 4)

OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) 10x-35=33-18x-12 10x+18x=33-12+35 28x=56 x=56/28 x=2 ans

OpenStudy (anonymous):

Proof: 5(2*2-7)=33-3(6*2+4) 5(4-7)=33-3(12+4) 5*-3=33-3*16 -15=33-48 -15=-15

OpenStudy (swag):

Thanks

OpenStudy (anonymous):

: )